# WaterMax: breaking the LLM watermark detectability-robustness-quality trade-off

### v1

@@ Appendix F Further experimental results
Every watermark algorithm is applied on a LLM working at temperature $1.0$ with nucleus sampling using the top $95\%$ token probabilities. The maximum text size is fixed at $1024$ tokens. The window size is fixed at $4$ for every algorithm. For both the Distribution shift and Exponential algorithms, we use the implementation from (Fernandez et al., 2023). Our code to replicate the benchmark is available on Github.
F.1 Quality
We herein provide further results regarding the impact of WaterMax on text quality compared to the state-of-the-art. We show in Figure 11 the perplexity of each watermark scheme as a function of their parameters. Observe that perplexity seems to stay constant for WaterMax contrary to other algorithms. In practice, we could not find a practical $n$ for which perplexity increased significantly. On the other hand, it does increases for increasing $N$, albeit extremely slowly compared to other algorithms (see the zoomed in figure).
In Figure 12 we do a similar experiment but using the MAUVE metric (Pillutla et al., 2021). The same conclusions as for perplexity can be made except this time we never observe a significant deviation from $1$ when using WaterMax. Since MAUVE is a lot noiser than perplexity, we favored the latter for comparing with other schemes.
F.2 Detectability
We herein provide further results regarding the detectability of WaterMax compared to the state-of-the-art.
(a) Llama-2-7b-chat-hf
(b) Mistral-7B-Instruct-v0.2
Figure 13: (log) Median false-positive rate as a function of relative perplexity for different watermarking schemes. Note that the relative perplexity for the Exponential scheme can get lower than $1$ due to the temperature of the watermark generator being lower than the temperature used to generate the original text.
In order to provide a better understanding of the trade-off between quality and detectability for each algorithm, we show in Figure 13 the false-postive rate of the detector for a power of $50\%$ as a function of the perplexity using OPT-2.7b as an oracle. We show the results for two different LLMs: Llama-2-7b-chat and Mistral-7b-Instruct-v0.2. If we look at Llama-2’s results, we observe that WaterMax allows to obtain far better false-positive rates with a minimal impact on quality, with the other algorithms struggling to reach these performances except by substantially increasing the perplexity. Note that the Exponential scheme can go below a relative perplexity of $1$ since it directly modify the temperature of the LLM to tune its strength. This can give it an “unfair” advantage since we could similarly improve the performance of both our scheme and the Distribution shift by increasing the temperature which would increase the entropy of the token distribution. Despite this advantage, the Exponential scheme performance stays far below both schemes. However, if we look at the performance for Mistral, this is reversed. Now the Exponential scheme can attain an extremely low false-positive rate even with a temperature close to $1$. Our scheme still largely outperforms Distribution shift for quality loss close to $0\%$ but necessitates the use of a high number of chunks to equalize the performance of the Exponential scheme.
This shows the high impact of the LLM on watermark schemes: the three schemes are dependent on the entropy of token distributions to perform well. The Distribution shift and Exponential schemes are extremely sensitive to low entropy distributions since they rely watermark “token by token”, hence the large variations of performance between Llama-2 (lower entropy model) and Mistral (higher entropy model) for the Exponential scheme. On the other hand, WaterMax is far more resilient since it better utilizes the entropy by working on chunks of tokens. As an example, the false positive rate of the Exponential scheme is almost divided by two between Llama-2 and Mistral at temperature $1$ whereas our scheme only looses approximately $20\%$ at $N=64$. Note that the higher the number of chunks $N$, the fewer tokens per chunks and hence the less efficient our scheme becomes.
In Figure 14-16 we provide the false-positive rate of the detector for a power of $50\%$ as a function of the size of the text. This is close to the ‘watermark size” defined in (Piet et al., 2023) but allows a better overview of the dependency of the watermarks on the text size.
F.3 Robustness
In Section 7 of this paper, we followed (Piet et al., 2023) regarding the measurement of the robustness of a watermark. Even though their method allows to summarize robustness under a single value, it does not tell the whole story since it is computed only for a specific false-alarm rate ($2\%$ in their case) and because it does not take into account the original quality of the watermarked text. Indeed, a direct consequence of the tamper-resistance as defined in (Piet et al., 2023) is that one can make a watermark more tamper-resistant by simply increasing its strength. It thus makes little sense to compare algorithms’ tamper-resistance at different quality regimes. Furthermore we found that some attacks that receives a high rating by Llama-2 actually substantially alter the readability of the text. In particular, the Typos-0.1 attacks, which randomly inserts typos in almost every word in the text, makes the resulting text barely legible, yet the attacked texts still obtain a Llama-2 rating within $90\%$ of the of the original watermarked text.
For all these reasons, we report results for individual attacks and for schemes working at a similar relative perplexity of $1.10$ in Figure 17-18. Here the story is quite different than the one told using the tamper-resistance metric. If we look at the absolute performance of the watermarks, they do not differ from the results obtained on detectability with no attacks: Distribution Shift performs the worse at this perplexity regime whereas the WaterMax ($N=64$) and Exponential schemes rankings depend on the LLM used. On the other hand, if we look at the watermark performance relative to the case where not attack is performed, every watermark seems to perform approximately the same.
This can be explained by leveraging the model presented in Section 5. Recall that any attack has the same impact on a watermarked text: changing one token modifies up to $h$ $p$-values such that they are distributed under $\mathcal{H}_{0}$ instead of $\mathcal{H}_{1}$. Under this model, we can simulate any attack by randomly modifying $(1-\alpha)\%,0\leq\alpha\leq 1$ of the scores. We can then plot the relative loss of power of the detector as a function of $\alpha$ to study the robustness of the scheme to attacks. Note that this analysis is general in the sense that it does not depend on the existence of an attack at a given $\alpha$ which still results in a text with sufficient quality. We believe the role of the watermark designer is to ensure the highest possible detectability of watermark whatever the strength of the attack; the design of attacks preserving text quality is the burden of the attacker. An analysis of the robustness of a watermark should thus be comprehensive, providing the performance of the watermark under any possible type of attacks, existing or not, and independent of text quality. We plot the average curves for each watermark in Figure 19. As can be observed, both WaterMax and Distribution Shift have identical robustness for $N=64$. WaterMax gets slightly more robust for low $(1-\alpha)$ by lowering $N$. However, under this analysis, the Exponential scheme is shown to be the most robust scheme, slightly outperforming both WaterMax and Distribution Shift for both Llama-2 and Mistral.
(a) Llama-2-7b-chat-hf
(b) Mistral-7B-Instruct-v0.2
Figure 14: Median false-positive rate as a function of text size in token for the WaterMax scheme for $N=64$. Due to the text size not necessarily reaching $1024$ tokens, the average “effective” $N$ is actually equal to $41$.
(a) Llama-2-7b-chat-hf
(b) Mistral-7B-Instruct-v0.2
Figure 15: Median false-positive rate as a function of text size in token for the Distribution Shift scheme.
(a) Llama-2-7b-chat-hf
(b) Mistral-7B-Instruct-v0.2
Figure 16: Median false-positive rate as a function of text size in token for the Exponential scheme.
(a) Llama-2-7b-chat-hf
(b) Mistral-7B-Instruct-v0.2
Figure 17: Median false-positive rate as a function of attack on the text. The results are given for watermark parameters leading to a relative perplexity of $1.10$.
(a) Llama-2-7b-chat-hf
(b) Mistral-7B-Instruct-v0.2
Figure 18: Median false-positive rate as a function of attack on the textrelative to the Median false-positive rate when no attack is performed . The results are given for watermark parameters leading to a relative perplexity of $1.10$.
(a) Llama-2-7b-chat-hf
(b) Mistral-7B-Instruct-v0.2
Figure 19: Median false-positive rate as a function of the ratio of modified token scores ($1-\alpha$). For each watermarking scheme, the curves are averaged for all studied parameters in Figure 14-16.

### v2

@@ Appendix I Impact of hashing window size and beam-search tokens
All texts are generated using Llama-2-chat-hf with nucleus sampling ($top_{p}=0.95$). The hashing window is fixed to $h=4$ for all algorithms. WaterMax is used with $b=0$ (no beam-search), which entails a virtually lossless scheme. Every watermarking scheme, except for Aaronson’s, is applied on a LLM working at temperature $\theta=1.0$.
K.1 Quality
Figures 15(a)-16(c) report the relative perplexity using opt-2.7b as an oracle as well as the ROUGE-L score for Aaronson, KGW, and WaterMax as a function of their respective parameter.
(a) KGW
(b) Aaronson
(c) Watermax $n=14,b=0$
(d) Focus on Figure 15(c). Note here that the relative perplexity of WaterMax does not exceed $1.2$.
Figure 15: Relative perplexity of watermarked text over non-watermarked text measured using opt-2.7b as an oracle.
(a) KGW
(b) Aaronson
(c) Watermax $n=14,b=0$
Figure 16: ROUGE-L
K.2 Impact of the text size
We generate texts of a maximum size of $L=1024$ tokens. For such long texts, the detectability easily reaches $P_{D}\approx 1.0$ if fixing $P_{FA}=10^{-6}$. Instead, we fix a detectability of $P_{D}=0.5$ and compute the corresponding $p$-values. This metric is close to the definition of ‘watermark size’ defined by Piet et al. [28]. The advantage of this metric is that we can report an increase in watermark performance even for long texts as $p$-svalues tend to decrease substantially with text size.
(a) KGW
(b) Aaronson
(c) Watermax ($b=0$) with chunk size fixed to $l=16$ tokens. Due to text size not always reaching $1024$ tokens, the effective $N$ is closer to $41$.